A ball is thrown horizontally from the top of a 103-m building for 4.58 seconds. What are the horizontal and vertical velocities of the ball if its initial velocity is 3 m/s?


1
Expert's answer
2022-04-29T15:30:01-0400

the horizontal velosity is constant 3 m/s

the vertical velocity at 4.58 s is given by

"v=gt^2/2=9.8*4.58^2/2=103\:\rm m"